# Formatting Instructions For NeurIPS 2024

### v1

@@ 4.2.2 General graphs
Given the underlying graph structure, we can construct the linear map $H$ from Equation (16) and the analytical solution for $C_{\theta\phi}$ in Appendix B. Finding the set of eigenvalues is then a standard task. We show the applicability of our formalism in Figure 2 by showing that GCNs degenerate to a zero distance state state exactly when $\sigma_{w}^{2}<\sigma_{w,\mathrm{crit}}^{2}$. Panel a) shows the maximal eigenvalue of the linearized GCN GP dynamics around the zero distance state, $\max\{\lambda^{p}\}$, together with the maximum feature distance between any two pairs of nodes in the graph, $\max_{(\alpha,\beta)}\{d(\boldsymbol{x}_{\alpha},\boldsymbol{x}_{\beta})\}$. The maximum feature distance increases from zero at the point where the state $K_{\alpha\beta}^{(l)}=k^{(l)}$ becomes unstable. This means that beyond this point, the GCN has feature vectors that differ across nodes and therefore does not oversmooth. This is explicitly shown in panels b) and c), where the equilibrium feature distance is plotted as a heatmap. At point $A$ (panel b)), within the oversmoothing phase, all equilibrium feature distances are indeed zero, the network therefore converged to a state in which all features are the same. At point $B$ (panel c)) on the other hand, pairs of nodes exist that have finite distance. Panel d) shows the feature distance averaged within the same community and across communities as a function of layers. While both converge to finite values, the averaged feature distance is larger for pairs of nodes across communities. The equilibrium state thus contains information about the graph topology. This aspect can also be observed in the block structure of the distance matrix in panel c). We thus showed that also on more general graphs the presented formalism predicts the transition point between the oversmoothing and the non-oversmoothing phase, corresponding to a transition between regular and chaotic behavior.
@@
Figure 2: The non-oversmoothing phase in a contextual stochastic block model instance with parameters $N=100$, $d=5$, $\lambda=1$. The shift operator is chosen according to (6) with $g=0.3$, and $\sigma_{b}^{2}=0$ and $\phi(x)=\mathrm{erf}(\frac{\sqrt{\pi}}{2}x)$. a) The maximum feature distance between any pair of two nodes in equilibrium (blue) and the largest eigenvalue of the linearized GCN GP dynamics at the zero distance state as a function of weight variance $\sigma_{w}^{2}$. The red line marks the point where $\max_{i}\{|\lambda_{i}^{\mathrm{p}}|\}=1$. b) Heatmap of the equilibrium distance matrix with entries $d_{\alpha\beta}=d(\boldsymbol{x}_{\alpha},\boldsymbol{x}_{\beta})$ (Equation (13)) at $\sigma_{w}^{2}=1.3$, marked as point $A$ in panel a). Colorbar shared with the plot in c). c) Same as b) but at point $B$ with $\sigma_{w}^{2}=2$. d) Features distances $d_{\alpha\beta}^{(l)}=d(\boldsymbol{x}_{\alpha}^{(l)},\boldsymbol{x}_{\beta}^{(l)})$ as a function of layers for random inputs $X_{\alpha i}^{(0)}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,1)$ and a finite-size GCN with $d_{l}=200$, averaged for distances for pairs of nodes within the same community (red) and across communities (purple).
@@
We discuss how the assumptions on weight matrices in related theoretical work [2, 40] exclude networks in the chaotic phase in Appendix E, explaining why the non-oversmoothing phase has not been reported before.
@@ 4.3 Implications for performance
Lastly we want to investigate the implications of the non-oversmoothing phase on performance. We do this by applying the GCN GP as well as a finite-size GCN to the task of node classification in the CSBM model and measure their performance, shown in Figure (3). Panel a) shows the generalization error (mean squared error) averaged over multiple instances of CSBMs dependent on the weight variance $\sigma_{w}^{2}$ and number of layers $l$. In the oversmoothing phase where most GCNs in the literature are initialized (see Appendix E), the generalization error increases significantly already after only a couple of layers. We observe the best performance near the transition to chaos where the GCN GP stays informative up to $100$ layers. Panel b) shows the performance of the GCN GP as a function of $\sigma_{w}^{2}$ for a fixed number of layers ($L=1,4,16,64,256,1024$ from turquoise to dark blue). While the generalization error increases to one (being random chance) in the oversmoothing phase, GCN GPs in the chaotic phase stay informative even at more than a thousand layers. This can be explained by Figure 2: For such deep networks, the dynamics are very close to the equilibrium and thus no information of the input features $\boldsymbol{X}^{(0)}$ is transferred to the output. The state, however, still contains information of the network topology from the adjacency matrix, leading to better than random chance performance. In panel c) we explicitly show the layer dependence of the generalization error for the GCN GP at the critical point ($\sigma_{w}^{2}=\sigma_{w,\mathrm{crit}}^{2}$), in the oversmoothing ($\sigma_{w}^{2}=\sigma_{w,\mathrm{crit}}^{2}-1$) and in the chaotic phase ($\sigma_{w}^{2}=\sigma_{w,\mathrm{crit}}^{2}+1$). Again, we see a fast performance drop for oversmoothing networks, while in the chaotic phase and at the critical point the GCN GP obtain good performance also at large depths, whereby performance peaks at $L\approx 15$ layers. Thus tuning the weight variance does not only prevent oversmoothing, but may also come with computational benefits by allowing additional layers.
@@
Figure 3: Generalization error (mean squared error) of the Gaussian process for a CSBM with parameters $N=20$, $d=5$, $\lambda=1$, $\gamma=1$ and $\mu=4$. The shift operator is defined in (6) with $g=0.9$, other parameters are $\sigma_{b}^{2}=0$, $\phi(x)=\mathrm{erf}(\frac{\sqrt{\pi}}{2}x)$ and $\sigma_{ro}=0.01$. In both panels we use $N^{\mathrm{train}}=10$ training nodes and $N^{\mathrm{test}}=10$ test nodes, five training nodes from each of the two communities. Labels are ±$\pm 1$ for the two communities, respectively. For all panels, we show averages over $50$ CSBM instances. a) Heatmap of the generalization error of the GCN GP dependent on number of layers $L$ and weight variance $\sigma_{w}^{2}$. The red line shows the transition to the non-oversmoothing phase. b) Generalization error dependent on weight variance $\sigma_{w}^{2}$ and $L=1,4,16,64,256,1024$ from turquoise to dark blue. c) Generalization error dependent on the layer for the GCN GP at the critical line $\sigma_{w}^{2}=\sigma_{w,\mathrm{crit}}^{2}$, in the oversmoothing phase $\sigma_{w}^{2}=\sigma_{w,\mathrm{crit}}^{2}-1$ and the non-oversmoothing phase $\sigma_{w}^{2}=\sigma_{w,\mathrm{crit}}^{2}+1$. d) Performance of randomly initialized finite-size GCNs with $d_{l}=200$ for $l=1,\dots,L$ where only the linear readout layer is trained with gradient descent (details in Appendix D) at the critical line $\sigma_{w}^{2}=\sigma_{w,\mathrm{crit}}^{2}$, in the oversmoothing phase $\sigma_{w}^{2}=\sigma_{w,\mathrm{crit}}^{2}-1$ and the non-oversmoothing phase $\sigma_{w}^{2}=\sigma_{w,\mathrm{crit}}^{2}+1$.
@@ 5 Discussion
In this study we used the equivalence of GCNs and GPs to investigate oversmoothing, the property that features at different nodes converge over layers to the same feature vector in an exponential manner. By extending concepts such as the propagation depth and chaos from the study of conventional deep feedforward neural networks [34], we are able to derive a condition to avoid oversmoothing. This condition is phrased in terms of an eigenvalue problem of the linearized GCN GP dynamics around the state where all features are the same: This state is stable if all eigenvalues are smaller than one, thus the networks do oversmooth. If one eigenvalue, however, is larger than one, the state where the features are the same on all nodes becomes unstable. While most GCNs studied in the literature are in the oversmoothing phase [2, 40], the non-oversmoothing phase can be reached by a simple tuning of the weight variance at initialization. An analogy can be drawn between the chaotic phase of DNNs and the non-oversmoothing phase of GCNs. Previous theoretical works have proven that oversmoothing is inevitable in some GNN architectures, among them GCNs; these works, however, make crucial assumptions on the weight matrices, constraining their variances to be in what we identify as the oversmoothing phase.
@@
Limitations. The current analysis is based on the equivalence of GCNs and GPs which strictly holds only in the limit of infinite feature dimension. GCNs with large feature vectors ($d_{l}=200$) are well described by the theory, as shown in Section 4. For a small number of feature dimensions, however, we expect deviations from the asymptotic results. Throughout this work, we assumed a column-stochastic shift operator which made the equilibrium $K^{\mathrm{eq}}$ in the oversmoothing phase particularly simple. For other shift operators, we expect the equilibrium $K^{\mathrm{eq}}$ to look different in detail. Yet, we hypothesize that increasing the weight variance makes the equilibrium more informative of the graph topology, as in the stochastic shift operator case. The choice of nonlinearity is unrestricted, but in the general case numerical integration of (8) is needed.
@@
To determine whether a given weight variance is in the non-oversmoothing phase, one calculates the eigenvalues of the linearized GCN GP dynamics which take the form of an $N^{2}\times N^{2}$ matrix (see Equation (16)), this has a run time of $\mathcal{O}(N^{6})$. While this becomes computationally expensive for large graphs, the conceptual insights of the presented analysis remain. In practical applications with large graphs one may reduce the computational load by determining the transition point via computation of the GCN GP prior until it is close to equilibrium. This procedure has a runtime of $\mathcal{O}(N^{2}L^{\mathrm{eq}})$ where $L^{\mathrm{eq}}$ is the number of layers after which the process is sufficiently close to equilibrium. One might then do an interval search on the weight variance until the transition point is determined with sufficient accuracy.
@@
Oulook. Formulating GCNs with the help of GPs can be considered the leading order in the number of feature space dimension $d_{l}$ when approximating finite-size GCNs. Computing corrections for finite numbers of hidden feature dimensions would allow the characterization of feature learning in such networks, similar as in standard deep networks [23, 43, 35]. Moreover, the generalization of this formalism to more general shift operators and other GNN architectures [44, 4] are possible directions of future research.
@@ Acknowledgements
Appendix A The linearized GP of GCNs
We start from the full GCN GP iterative map which we restate here for readability
$K_{\alpha\beta}^{(l+1)}=$ $T_{\alpha\beta}[K^{(l)}]=\sigma_{b}^{2}+\sigma_{w}^{2}\,\sum_{\gamma,\delta}A_{\alpha\gamma}A_{\beta\delta}\,\Big{\langle}\phi\Big{(}h_{\gamma}^{(l)}\Big{)}\phi\Big{(}h_{\delta}^{(l)}\Big{)}\Big{\rangle}_{h_{\gamma}^{(l)},h_{\delta}^{(l)}}$ (21)
where $h_{\gamma}^{(l)}$ and $h_{\delta}^{(l)}$ are drawn from a 00-mean Gaussian distribution with covariance $\langle h_{\gamma}^{(l)}h_{\delta}^{(l)}\rangle=K_{\gamma\delta}^{(l)}$. The full covariance matrix of layer $l$ is denoted as $K^{(l)}$. Here, we differentiate between the iterative maps $T_{\alpha\beta}:\mathbb{R}^{N\times N}\to\mathbb{R}$ of which there are $N^{2}$, one for each pair of nodes $\alpha$ and $\beta$, and the entries of the covariance matrix in the layer $l+1$ being $K_{\alpha\beta}^{(l+1)}$. Notice that $T_{\alpha\beta}=T_{\beta\alpha}$ due to the symmetry of the covariance matrix. In the maps $T_{\alpha\beta}$, the covariance matrix of the previous layer only shows up in the expectation value $C_{\gamma\delta}=\Big{\langle}\phi\Big{(}h_{\gamma}^{(l)}\Big{)}\phi\Big{(}h_{\delta}^{(l)}\Big{)}\Big{\rangle}_{h_{\gamma}^{(l)},h_{\delta}^{(l)}}$ such that the linearized dynamics around a fixed point (being equilibrium $K_{\alpha\beta}^{\mathrm{fix}}=K_{\alpha\beta}^{\mathrm{eq}}$ or zero distance state $K_{\alpha\beta}^{\mathrm{fix}}=k$) with $K_{\alpha\beta}^{(l)}=K_{\alpha\beta}^{\mathrm{fix}}+\Delta_{\alpha\beta}^{(l)}$ read

### v2

@@ 4.2.2 General graphs
Figure 2: The non-oversmoothing phase in a contextual stochastic block model instance with parameters $N=100$, $d=5$, $\lambda=1$. The shift operator is chosen according to (6) with $g=0.3$, and $\sigma_{b}^{2}=0$ and $\phi(x)=\mathrm{erf}(\frac{\sqrt{\pi}}{2}x)$. a) The maximum feature distance between any pair of nodes in equilibrium obtained from computing the GCN GP prior for $L=4,000$ layers (blue) and the largest eigenvalue of the linearized GCN GP dynamics at the zero distance state as a function of weight variance $\sigma_{w}^{2}$. The red line marks the point where $\max_{i}\{|\lambda_{i}^{\mathrm{p}}|\}=1$. b) Heatmap of the equilibrium distance matrix with entries $d_{\alpha\beta}=d(\boldsymbol{x}_{\alpha},\boldsymbol{x}_{\beta})$ (Equation (13)) at $\sigma_{w}^{2}=1.3$, marked as point $A$ in panel a). Colorbar shared with the plot in c). c) Same as b) but at point $B$ with $\sigma_{w}^{2}=2$. d) Features distances $d_{\alpha\beta}^{(l)}=d(\boldsymbol{x}_{\alpha}^{(l)},\boldsymbol{x}_{\beta}^{(l)})$ as a function of layers for random inputs $X_{\alpha i}^{(0)}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,1)$ and a finite-size GCN with $d_{l}=200$, averaged for distances for pairs of nodes within the same community (red) and across communities (purple).
Given the underlying graph structure, we can construct the linear map $H$ from Equation (16) and the analytical solution for $C_{\theta\phi}$ in Appendix A. Finding the set of eigenvalues is then a standard task. We show the applicability of our formalism in Figure 2 by showing that GCNs degenerate to a zero distance state state exactly when $\sigma_{w}^{2}<\sigma_{w,\mathrm{crit}}^{2}$. Panel a) shows how this procedure correctly predicts the transition in the given CSBM instance: The maximum feature distance between any pair of nodes increases from zero at the point where the state $K_{\alpha\beta}^{(l)}=k^{(l)}$ becomes unstable. This means that beyond this point, the GCN has feature vectors that differ across nodes and therefore does not oversmooth. This is more explicitly shown in panels b) and c), where the equilibrium feature distance is plotted as a heatmap. At point $A$ (panel b)), within the oversmoothing phase, all equilibrium feature distances are indeed zero, the network therefore converges to a state in which all features are the same. At point $B$ (panel c)) on the other hand, pairs of nodes exist that have finite distance. In the latter case, one can recognize the community structure of the CSBM: the lower left and upper right quadrants are lighter than the diagonal ones, indicating larger feature distances across communities than within. The equilibrium state thus contains information about the graph topology. This phenomenon is also observed in panel d), where we show the predicted feature distance averaged for nodes within or between classes as a function of layers compared to finite-size simulations. Again, theoretical predictions match with simulations. Thus also on more general graphs the presented formalism predicts the transition point between the oversmoothing and the non-oversmoothing phase, corresponding to a transition between regular and chaotic behavior.
@@
We discuss how the assumptions on weight matrices in related theoretical work [2, 45] exclude networks in the chaotic phase in Appendix D, explaining why the non-oversmoothing phase has not been reported before. In Appendix E we observe how increasing the weight variance increases the oversmoothing measure $\mu(X)$ in equilibrium also in the case of the more common shift operator proposed in the original work [19], despite the fact that this shift operator does not have the oversmoothed fixed point in the sense of Equation (19).
@@ 4.3 Implications for performance
Lastly we want to investigate the implications of the non-oversmoothing phase on performance. We do this by applying the GCN GP as well as a finite-size GCN to the task of node classification in the CSBM model and measure their performance, shown in Figure 3. Panel a) shows how the generalization error of the GCN GP changes depending on the weight variance $\sigma_{w}^{2}$ and the number of layers $L$. In the oversmoothing phase where most GCNs in the literature are initialized (see Appendix D), the generalization error increases significantly already after only a couple of layers. We observe the best performance near the transition to chaos where the GCN GP stays informative up to $100$ layers. In panel b) we test the generalization error for even deeper networks. While the generalization error increases to one (being random chance) in the oversmoothing phase, GCN GPs in the chaotic phase stay informative even at more than a thousand layers. This can be explained by Figure 2: For such deep networks, the dynamics are very close to the equilibrium and thus no information of the input features $\boldsymbol{X}^{(0)}$ is transferred to the output. The state, however, still contains information of the network topology from the adjacency matrix, leading to better than random chance performance. In panel c) we explicitly show the layer dependence of the generalization error for the GCN GP at the critical point, in the oversmoothing and in the chaotic phase. Again, we see a fast performance drop for oversmoothing networks, while in the chaotic phase and at the critical point the GCN GP obtains good performance also at large depths, with performance peaking at $L\approx 15$ layers. Tuning the weight variance thus not only prevents oversmoothing, but may also allow the construction of GCNs with more layers and possibly better generalization performance.
Figure 3: Generalization error (mean squared error) of the Gaussian process for a CSBM with parameters $N=20$, $d=5$, $\lambda=1$, $\gamma=1$ and $\mu=4$. The shift operator is defined in (6) with $g=0.1$, other parameters are $\sigma_{b}^{2}=0$, $\phi(x)=\mathrm{erf}(\frac{\sqrt{\pi}}{2}x)$ and $\sigma_{ro}=0.01$. In all panels we use $N^{\mathrm{train}}=10$ training nodes and $N^{\mathrm{test}}=10$ test nodes, five training nodes from each of the two communities. Labels are ±$\pm 1$ for the two communities, respectively. For all panels, we show averages over $50$ CSBM instances. a) Heatmap of the generalization error of the GCN GP dependent on number of layers $L$ and weight variance $\sigma_{w}^{2}$. The red line shows the transition to the non-oversmoothing phase. b) Generalization error dependent on weight variance $\sigma_{w}^{2}$ and depths $L=1,4,16,64,256,1024$ from turquoise to dark blue. c) Generalization error dependent on the layer for the GCN GP at the critical line $\sigma_{w}^{2}=\sigma_{w,\mathrm{crit}}^{2}$, in the oversmoothing phase $\sigma_{w}^{2}=\sigma_{w,\mathrm{crit}}^{2}-1$ and the non-oversmoothing phase $\sigma_{w}^{2}=\sigma_{w,\mathrm{crit}}^{2}+1$. d) Performance of randomly initialized finite-size GCNs with $d_{l}=200$ for $l=1,\dots,L$ where only the linear readout layer is trained with gradient descent (details in Appendix F) at the critical line $\sigma_{w}^{2}=\sigma_{w,\mathrm{crit}}^{2}$, in the oversmoothing phase $\sigma_{w}^{2}=\sigma_{w,\mathrm{crit}}^{2}-1$ and the non-oversmoothing phase $\sigma_{w}^{2}=\sigma_{w,\mathrm{crit}}^{2}+1$.
@@
Figure 4: GCN GP performance on the Cora datset [38]. a) Generalization error (mean squared error) as a function of layers $L$ and weight variance $\sigma_{w}^{2}-\sigma_{w,\text{crit.}}^{2}$ for our stochastic shift operator (6) with $g=0.9$. The value of $\sigma_{w,\text{crit.}}^{2}\approx 1$ is determined numerically in Appendix G. b) Layer dependent generalization error and accuracy for GCNs near the transition $\sigma_{w}^{2}=\sigma_{w,\text{crit.}}^{2}+0.1$. Grey dashed line shows accuracy obtained for GCNs in the original work [19]. Numerical details in Appendix F.
Additionally, we test the performance of the GCN GP on the real world citation network Cora [38]. Evaluating the eigenvalue condition (20) would be computationally expensive for such a large dataset, therefore we find the transition by numerically evaluating the feature distance $\mu(X)$ in equilibrium and search for the $\sigma_{w}^{2}$ at which this distance becomes non-zero. This procedure results in $\sigma_{w,\text{crit}}^{2}\approx 1$ and is presented in more detail in Appendix G. Figure 4 panel a) shows the performance of the GCN GP dependent on the number of layers $L$ and weight variance $\sigma_{w}^{2}$: as for the CSBM in Figure 3 we observe that the performance for deep GCN GPs is best near the transition in the non-oversmoothing phase. Furthermore, GCN GPs with more layers achieve lower generalization error. This is shown more directly in panel b). There we observe the layer dependence for GCN GPs near the transition in the non-oversmoothing regime. Indeed, the accuracy increases up to a hundred layers, reaching the accuracy of finite-size GCNs stated in [19].
Near the transition, accuracy increases for up to $L=100$, and the generalization error improves even beyond this. We hypothesize that this many layers are required for high performance partly due to our choice of the shift operator. The Cora dataset has a maximum degree $d_{\text{max}}=168$ leading to small off-diagonal elements for the choice of our shift operator: Recall that in Equation (6), the parameter $g$ is constrained to be $g\in(0,1)$. As a consequence, the off-diagonal elements of the shift operator are $A_{ij}<\frac{1}{d_{\text{max}}}=\frac{1}{168}$. Many convolutional layers are then needed to incorporate information from a node’s neighbors.
One might wonder whether it is possible to initialize weights in say the oversmoothing regime, and transition to the non-oversmoothing regime during training. We argue that this is possible in the case of Langevin training (Appendix H).
@@ 5 Discussion
In this study we used the equivalence of GCNs and GPs to investigate oversmoothing, the property that features at different nodes converge over layers to the same feature vector in an exponential manner. By extending concepts such as the propagation depth and chaos from the study of conventional deep feedforward neural networks [37], we are able to derive a condition to avoid oversmoothing. This condition is phrased in terms of an eigenvalue problem of the linearized GCN GP dynamics around the state where all features are the same: This state is stable if all eigenvalues are smaller than one, thus the networks do oversmooth. If one eigenvalue, however, is larger than one, the state where the features are the same on all nodes becomes unstable. While most GCNs studied in the literature are in the oversmoothing phase [2, 45], the non-oversmoothing phase can be reached by a simple tuning of the weight variance at initialization. An analogy can be drawn between the chaotic phase of DNNs and the non-oversmoothing phase of GCNs. Previous theoretical works have proven that oversmoothing is inevitable in some GNN architectures, among them GCNs; these works, however, make crucial assumptions on the weight matrices, constraining their variances to be in what we identify as the oversmoothing phase. Near the transition, we find GCNs which are both deep and expressive, matching the originally reported GCN performance [19] on the Cora dataset with GCN GPs beyond $100$ layers.
@@
Limitations. The current analysis is based on the equivalence of GCNs and GPs which strictly holds only in the limit of infinite feature dimension. GCNs with large feature vectors ($d_{l}=200$) are well described by the theory, as shown in Section 4. For a small number of feature dimensions, however, we expect deviations from the asymptotic results. Throughout the main part of this work, we assumed a row-stochastic shift operator which made the equilibrium $K^{\mathrm{eq}}$ in the oversmoothing phase particularly simple. For other shift operators, we expect qualitatively similar results while the equilibrium $K^{\mathrm{eq}}$ may look different in detail. In our preliminary experiments on the common shift operator from [19] (Appendix E), we indeed find that increasing the weight variance increases the distances between features also in this case. We hypothesize that this effect makes the equilibrium more informative of the graph topology, as in the stochastic shift operator case. The choice of non-linearity is unrestricted, but in the general case numerical integration of (8) is needed.
@@
To determine whether a given weight variance is in the non-oversmoothing phase, one calculates the eigenvalues of the linearized GCN GP dynamics which take the form of an $N^{2}\times N^{2}$ matrix (see Equation (16)), this has a run time of $\mathcal{O}(N^{6})$. While this becomes computationally expensive for large graphs, the conceptual insights of the presented analysis remain. In practical applications with large graphs one may reduce the computational load by determining the transition point via computation of the GCN GP prior until it is close to equilibrium. This procedure has a runtime of $\mathcal{O}(N^{3}L^{\mathrm{eq}})$ where $L^{\mathrm{eq}}$ is the number of layers after which the process is sufficiently close to equilibrium. One might then do an interval search on the weight variance until the transition point is determined with sufficient accuracy.
@@
Oulook. Formulating GCNs with the help of GPs can be considered the leading order in the number of feature space dimension $d_{l}$ when approximating finite-size GCNs. Computing corrections for finite numbers of hidden feature dimensions would allow the characterization of feature learning in such networks, similar as in standard deep networks [26, 48, 39]. Moreover, the generalization of this formalism to more general shift operators and other GNN architectures [49, 4] like GATs [41] are possible directions of future research. In the special case of GATs we expect similar results to the GCN analyzed here, since the shift operator is constructed using a softmax and therefore also is row-stochastic.
@@ Acknowledgements
Appendix A Analytical solution for expectation values
To evaluate our theory, we need to compute expectation values given in the form of (8) which we restate here for readability
where $h_{\gamma}^{(l)}$ and $h_{\delta}^{(l)}$ are zero mean random Gaussian variables with $\langle h_{\gamma}^{(l)}h_{\delta}^{(l)}\rangle=K_{\gamma\delta}^{(l)}$. This can be evaluated numerically for general non-linearities $\phi(x)$. For simplicity, however, we choose $\phi(x)=\mathrm{erf}(\frac{\sqrt{\pi}}{2}x)$ in our experiments where the scaling factor in the $\mathrm{erf}$ is chosen such that $\frac{\partial\phi}{\partial x}(0)=1$. In this case, the expectation value can be evaluated analytically to be
as shown in [44].
Appendix B The linearized GP of GCNs
We start from the full GCN GP iterative map (8,(9)), which we restate here for readability
$K_{\alpha\beta}^{(l+1)}=$ $T_{\alpha\beta}[K^{(l)}]=\sigma_{b}^{2}+\sigma_{w}^{2}\,\sum_{\gamma,\delta}A_{\alpha\gamma}A_{\beta\delta}C_{\gamma\delta}^{(l)}[K^{(l)}]$ (23)
where $h_{\gamma}^{(l)}$ and $h_{\delta}^{(l)}$ are drawn from a 00-mean Gaussian distribution with covariance $\langle h_{\gamma}^{(l)}h_{\delta}^{(l)}\rangle=K_{\gamma\delta}^{(l)}$. The full covariance matrix of layer $l$ is denoted as $K^{(l)}$. Here, we distinguish between the iterative maps $T_{\alpha\beta}:\mathbb{R}^{N\times N}\to\mathbb{R}$ of which there are $N^{2}$, one for each pair of nodes $\alpha$ and $\beta$, and the entries $K_{\alpha\beta}^{(l+1)}$ of the covariance matrix in the layer $l+1$. Notice that $T_{\alpha\beta}=T_{\beta\alpha}$ due to the symmetry of the covariance matrix. In the maps $T_{\alpha\beta}$, the covariance matrix of the previous layer only shows up in the expectation value $C_{\gamma\delta}=\Big{\langle}\phi\Big{(}h_{\gamma}^{(l)}\Big{)}\phi\Big{(}h_{\delta}^{(l)}\Big{)}\Big{\rangle}_{h_{\gamma}^{(l)},h_{\delta}^{(l)}}$ such that the linearized dynamics around a fixed point (being equilibrium $K_{\alpha\beta}^{\mathrm{fix}}=K_{\alpha\beta}^{\mathrm{eq}}$ or zero distance state $K_{\alpha\beta}^{\mathrm{fix}}=k$) with $K_{\alpha\beta}^{(l)}=K_{\alpha\beta}^{\mathrm{fix}}+\Delta_{\alpha\beta}^{(l)}$ read